OpenStudy (xxlyh):

What is the value of z in the equation 5(2z + 4) = 0? -2 0 2 20

OpenStudy (xxlyh):

@Joshtray

Mehek (mehek14):

distribute first 5*2z = ? 5*4 = ?

OpenStudy (xxlyh):

20 and 10?

Mehek (mehek14):

10z + 20 = 0 subtract 20 0 - 20

OpenStudy (xxlyh):

subtract 20 from 0?

Mehek (mehek14):

ye 0 - 20

OpenStudy (richardaperez):

yes to get -20 on that side

OpenStudy (xxlyh):

-20

Mehek (mehek14):

-20 / 10

OpenStudy (xxlyh):

-2

OpenStudy (richardaperez):

yes

Mehek (mehek14):

that's your answer

OpenStudy (xxlyh):

oh lool thank you!!

Mehek (mehek14):

yw :)

OpenStudy (xxlyh):

do you mind helping me with one more?

Mehek (mehek14):

i'll help

OpenStudy (xxlyh):

ok ill post a new question

OpenStudy (phi):

btw, often it's faster to first divide by the 5 , on both sides in this case you get 5(2z + 4) = 0 becomes 2x + 4 = 0 (because 0/5 is still 0) now do 2x= -4 x= -2